# MMInA: Benchmarking Multihop Multimodal Internet Agents

### v1

@@ B.2 More Related Works
Recent progress in multimodal learning, showcased by models like CLIP[34], DALL-E[35], Stable Diffusion[36], Flamingo[3], and GPT series, has led to significant improvements in areas such as zero-shot classification, image creation, and in-context learning. Although these models employ various algorithmic approaches, including contrastive learning, diffusion techniques, and auto-regressive models, they share a fundamental reliance on large datasets comprising image-text pairs. This commonality underscores the importance of such datasets in driving advancements in multimodal AI capabilities.
Webdataset [^1] https://github.com/webdataset/webdataset is a commonly used dataset as it contains thousands of image-text pairs data scraped from websites; LAINON-5b[39] is a dataset that contains 5.85 billion CLIP-filtered image-text pairs, where 2.32B contain English language; MIMIC-IT[19] is a dataset consists of 2.8 million multimodal instruction-response pairs equipped with rich in-context information, with 2.2 million unique instructions derived from images and videos; DataComp[10] is a newly brought dataset which consists of four stages: A) Deciding on a scale that fits within resource limitations. B) Creating a dataset, opting for either the filtering approach or the Bring Your Own Data (BYOD) track. C) Using a set architecture and specific hyperparameters to train a CLIP model on the created dataset. D) Assessing the performance of the trained model across a variety of downstream tasks.
@@
Instruction tuning is a common method used in LLMs training, which involves refining pre-trained LLMs using datasets formatted as instructions. This approach enhances the model’s ability to perform new, unseen tasks by simply following directions, thereby improving its zero-shot capabilities. Some notable models like ChatGPT[30], InstructGPT[31], FLAN[45, 6] are built on top of instruction tuning methods.
@@
, and the optimizing objective can be formulated as:
@@
1) Tranformation The efficacy of instruction tuning in the training of LMMs is significantly constrained by the limitations in length and type of data available in current Visual Question Answering (VQA) datasets. To address this, some researchers have opted to adapt the provided instructions, transforming the succinct answer data into extended sentences enriched with semantic details [51]. Other studies, such as in, reconstructed the answer by prompting ChatGPT to emulate the capabilities of advanced language models.
2) Self-Instruct LLaVA [23] extends the multimodal approach by converting images into descriptive texts and outlines of bounding boxes, then uses GPT-4 to create additional data within the context provided by initial examples.
@@
Agents designed for Graphical User Interfaces (GUIs) are crafted to streamline complex activities on digital devices like smartphones and desktops. These GUI agents may employ HTML as inputs, or alternatively, use screenshots to facilitate task execution in a broader context. Traditionally, research has revolved around training these agents in restrictive, static environments, a practice that deviates from human learning and hinders the agents’ ability to make decisions akin to humans. However, the emergence of large language models (LLMs) and large multimodal models (LMMs) equipped with vast web knowledge marks a pivotal shift towards achieving a more human-like intellect in agents, sparking a surge in research on LLM/LMM-enhanced autonomous agents. This section aims to explore the latest state-of-the-art (SOTA) developments in autonomous agents, examining both web GUI agents and mobile GUI agents.
@@
1) GUI Agents - Web Agents Despite the current progress of web agents discussed in the main paper, several works also explored the development of web agents. TravelPlanner [46] proposed a benchmark that provides a sandbox environment with tools for accessing nearly four million data records. It includes 1,225 planning intents and reference plans to evaluate the planning strategies of language agents by using tools; OmniACT [15] presented a dataset and benchmark for assessing an agent’s capability to generate executable programs for computer tasks. It uses the PyAutoGUI Python library to automate mouse and keyboard operations across different operating systems and web domains. It addresses the limitations of HTML-based agents by providing a multimodal challenge where visual cues are crucial, thus enabling a more robust understanding of UI elements, but it still showed the inability to handle native desktop applications or multi-application tasks; WEBLINX [26] also proposed a benchmark for conversational web navigation, addressing the problem of enabling a digital agent to control a web browser and follow user instructions in a multi-turn dialogue fashion. The method involves a retrieval-inspired model that prunes HTML pages by ranking relevant elements, addressing the issue of LLMs not being able to process entire web pages in real-time. The technology used includes a dense markup ranker for element selection and multimodal models that combine screenshots, action history, and textual website representation. The performance is evaluated on tasks like creating a task on Google Calendar, with the model’s ability to replicate human behavior when navigating the web; DUAL-VCR [16], leverages the “dual view” of HTML elements in webpage screenshots, contextualizing each element with its visual neighbors. This approach uses both textual and visual features to create more informative representations for decision-making.
@@
2) GUI Agents - Mobile Agents Besides web agents, mobile GUI agents are gaining more and more popularity, which are developed to handle intricate tasks automatically on digital devices like smartphones. ERICA [7] defines a system for interaction mining in Android applications, it employs a human-computer interaction approach to capture the data, making it scalable and capable of capturing a wide range of interactions. PIXEL [38] and Pix2Struct [18] showed promising capability in multilingual transfer and UI navigation respectively, but they struggle with language understanding tasks compared to text-only LMs like BERT, limiting their utility. Patch-and-Text Prediction (PTP) proposed in [11] leads to better language understanding capabilities by masking and recovering both image patches and text within screenshots; AppAgent [48] presented a multimodal agent that operates smartphone apps through low-level actions like tapping and swiping, mimicking human interactions. The agent learns app functionalities through exploration, either autonomously or by observing human demonstrations, and then applies this knowledge to execute tasks; Mobile-Agent [44], which uses visual perception tools to locate operations on a mobile device using screenshots. It involves OCR models for identifying visual and textual elements, while realizing self-planning and self-reflection to autonomously navigate mobile apps, with a benchmark called Mobile-Eval introduced for performance evaluation; SeeClick [5] is a visual GUI agent that operates solely on screenshots, bypassing the need for structured text. It employs Large Vision-Language Models (LVLMs) enhanced with GUI grounding pre-training to accurately locate screen elements based on instructions. The method involves automating the curation of GUI grounding data and creating a GUI grounding benchmark named ScreenSpot. It adapts universally to various GUI platforms and its reliance on screenshots. simplifying action space to clicking and typing.

### v2

@@ Appendix C More Related Works
Recent progress in multimodal learning, showcased by models like CLIPRadford et al. (2021), DALL-ERamesh et al. (2021), Stable DiffusionRombach et al. (2022), FlamingoAlayrac et al. (2022), and GPT series, has led to significant improvements in areas such as zero-shot classification, image creation, and in-context learning. Although these models employ various algorithmic approaches, including contrastive learning, diffusion techniques, and auto-regressive models, they share a fundamental reliance on large datasets comprising image-text pairs. This commonality underscores the importance of such datasets in driving advancements in multimodal AI capabilities.
Webdataset***https://github.com/webdataset/webdataset is a commonly used dataset as it contains thousands of image-text pairs data scraped from websites; LAION-5BSchuhmann et al. (2022) is a dataset that contains 5.85 billion CLIP-filtered image-text pairs, where 2.32B contain English language; MIMIC-ITLi et al. (2025) is a dataset consists of 2.8 million multimodal instruction-response pairs equipped with rich in-context information, with 2.2 million unique instructions derived from images and videos; DataCompGadre et al. (2023) is a newly brought dataset which consists of four stages: A) Deciding on a scale that fits within resource limitations. B) Creating a dataset, opting for either the filtering approach or the Bring Your Own Data (BYOD) track. C) Using a set architecture and specific hyperparameters to train a CLIP model on the created dataset. D) Assessing the performance of the trained model across a variety of downstream tasks.
@@
Instruction tuning is a common method used in LLM training, which involves refining pre-trained LLMs using datasets formatted as instructions. This approach enhances the model’s ability to perform new, unseen tasks by simply following directions, thereby improving its zero-shot capabilities. Some notable models like ChatGPTAchiam et al. (2023), InstructGPTOuyang et al. (2022), FLANWei et al. (2021); Chung et al. (2024) are built on top of instruction tuning methods.
@@
the optimizing objective can be formulated as:
@@
1) Transformation The efficacy of instruction tuning in the training of LMMs is significantly constrained by the limitations in length and type of data available in current Visual Question Answering (VQA) datasets. To address this, some researchers have opted to adapt the provided instructions, transforming the succinct answer data into extended sentences enriched with semantic details Zhao et al. (2023). Other studies, such as in, reconstructed the answer by prompting ChatGPT to emulate the capabilities of advanced language models.
2) Self-Instruct LLaVA Liu et al. (2023a) extends the multimodal approach by converting images into descriptive texts and outlines of bounding boxes, then uses GPT-4 to create additional data within the context provided by initial examples.
@@
Agents designed for Graphical User Interfaces (GUIs) are crafted to streamline complex activities on digital devices like smartphones and desktops. These GUI agents may employ HTML as inputs or use screenshots to facilitate task execution in a broader context. Traditionally, research has revolved around training these agents in restrictive, static environments, a practice that deviates from human learning and hinders the agents’ ability to make decisions akin to humans. However, the emergence of large language models (LLMs) and large multimodal models (LMMs) equipped with vast web knowledge marks a pivotal shift towards achieving a more human-like intellect in agents, sparking a surge in research on LLM/LMM-enhanced autonomous agents. This section aims to explore the latest state-of-the-art (SOTA) developments in autonomous agents, examining both web GUI agents and mobile GUI agents.
@@
1) GUI Agents - Web Agents Despite the current progress of web agents discussed in the main paper, several works also explored the development of web agents. TravelPlanner Xie et al. (2024) proposes a benchmark that provides a sandbox environment with tools for accessing nearly four million data records. It includes 1,225 planning intents and reference plans to evaluate the planning strategies of language agents by using tools; OmniACT Kapoor et al. (2024) presents a dataset and benchmark for assessing an agent’s capability to generate executable programs for computer tasks. It uses the PyAutoGUI Python library to automate mouse and keyboard operations across different operating systems and web domains. It addresses the limitations of HTML-based agents by providing a multimodal challenge where visual cues are crucial, thus enabling a more robust understanding of UI elements, but it still shows the inability to handle native desktop applications or multi-application tasks; WEBLINX Lu et al. (2024) also proposes a benchmark for conversational web navigation, addressing the problem of enabling a digital agent to control a web browser and follow user instructions in a multi-turn dialogue fashion. The method involves a retrieval-inspired model that prunes HTML pages by ranking relevant elements, addressing the issue of LLMs not being able to process entire web pages in real-time. The technology used includes a dense markup ranker for element selection and multimodal models that combine screenshots, action history, and textual website representation. The performance is evaluated on tasks like creating a task on Google Calendar, with the model’s ability to replicate human behavior when navigating the web; DUAL-VCR Kil et al. (2024) leverages the “dual view” of HTML elements in webpage screenshots, contextualizing each element with its visual neighbors. This approach uses both textual and visual features to create more informative representations for decision-making.
@@
2) GUI Agents - Mobile Agents Besides web agents, mobile GUI agents are gaining more and more popularity, which are developed to handle intricate tasks automatically on digital devices like smartphones. ERICA Deka et al. (2016) defines a system for interaction mining in Android applications. It employs a human-computer interaction approach to capture the data, making it scalable and capable of capturing a wide range of interactions. PIXEL Rust et al. (2023) and Pix2Struct Lee et al. (2023) show promising capability in multilingual transfer and UI navigation, respectively, but they struggle with language understanding tasks compared to text-only LMs like BERT, limiting their utility. Patch-and-Text Prediction (PTP) proposed in Gao et al. (2024) leads to better language understanding capabilities by masking and recovering both image patches and text within screenshots; AppAgent Zhang et al. (2025) presents a multimodal agent that operates smartphone apps through low-level actions like tapping and swiping, mimicking human interactions. The agent learns app functionalities through exploration, either autonomously or by observing human demonstrations, and then applies this knowledge to execute tasks; Mobile-Agent Wang et al. (2024), which uses visual perception tools to locate operations on a mobile device using screenshots. It involves OCR models for identifying visual and textual elements, while realizing self-planning and self-reflection to autonomously navigate mobile apps, with a benchmark called Mobile-Eval introduced for performance evaluation; SeeClick Cheng et al. (2024) is a visual GUI agent that operates solely on screenshots, bypassing the need for structured text. It employs Large Vision-Language Models (LVLMs) enhanced with GUI grounding pre-training to accurately locate screen elements based on instructions. The method involves automating the curation of GUI grounding data and creating a GUI grounding benchmark named ScreenSpot. It adapts universally to various GUI platforms and relies on screenshots. Simplifying the action space to clicking and typing.
C.1 Other Benchmarks
Shi et al. (2017) and Liu et al. (2018) establish a platform of website widgets where agents can complete online tasks through basic keyboard and mouse operations. APIBench Li et al. (2023b), introduced by GorillaPatil et al. (2024), is a tool-augmented LLM benchmark to assess the tool utilization abilities of agents for code generation tasks. AgentBench Liu et al. (2024) steps forward to provide a more general toolkit with lots of closed-box environments to assess agents’ performances in answering user queries.
Webshop Yao et al. (2022) builds a simulated e-commerce environment featuring 1.18 million real-world products, complemented by 12,087 crowdsourced textual instructions. It converted the action prediction task into a choice-based imitation learning process, which facilitated the accuracy and ability for task execution. However, this approach failed to evaluate open-ended agent actions in the real world. It was also limited by its monotonous design of a one-website environment, which resulted in only a single category of web browsing tasks. Mind2Web Deng et al. (2023) tries to construct a generalist web agent, which creates a dataset for crafting and benchmarking web agents by the ability of instruction following. It proposed a two-stage training to convert the action prediction problem into MCQs. SeeAct Zheng et al. (2024) is a following work that enabled multimodal information for visually understanding rendered webpages and generating more accurate action plans. WebVoyager He et al. (2024) is capable of capturing screenshots of web pages and then using JavaScript tools to automatically identify interactive elements based on the types of webpage elements. WebArena Zhou et al. (2024) deploys a standalone set of multicategory websites in an interactive environment. VisualWebArena Koh et al. (2024) is a subsequent project that built upon the foundation of WebArena, introducing the reliance on visual cues into the benchmark’s design. The tasks of existing benchmarks are oversimplified whose completions requiring a single website, which is highly diverged from the natural web browsing tasks and should originally be designed for multihop over a long-horizon setting.